# Efficient Multi Subject Visual Reconstruction from fMRI Using Aligned Representations

### v1

@@ 3.3 Reconstruction Pipeline
Our reconstruction pipeline follows a similar structure to previous works, dividing the reconstruction process into two components: a high-level semantic pipeline that maps to CLIP embeddings and a low-level visual pipeline that maps to visual autoencoder embeddings. Both pipelines receive the same fMRI data as input but are trained separately. The outputs from both pipelines are then integrated into a pre-trained diffusion model to generate the final reconstruction. The exact training process and loss functions are presented in Appendix Sections I and J.
@@
Flattened fMRI voxels from the nsdgeneral region are fed into a subject-specific adapter, consisting of a single linear layer with a GELU non-linearity. This maps the fMRI data into a subject-agnostic shared space. This shared-space representation is then processed by a Residual Multi-Layer Perceptron (MLP). A modality gap still exists between outputs from the MLP and the ground-truth CLIP (Radford et al., 2021) embedding space. We further incorporate a pre-trained diffusion prior based on DALL-E 2 (Ramesh et al., 2022) to bridge this modality gap. The output from the prior is the mapped CLIP embedding. This is fed as conditioning to the diffusion model.
Our high-level pipeline is based on the architecture used in MindEye1 (Scotti et al., 2023) and MindEye2 (Scotti et al., 2024), where the fMRI data is mapped to CLIP image embeddings of the corresponding source images. The pretraining phase leverages a combination of diffusion prior loss and CLIP loss. During fine-tuning, we add an additional MSE loss for the common images to further refine the mappings. The overall loss function in this stage is:
@@
$\mathcal{L}_{\text{new}}=\begin{cases}\lambda_{1}\mathcal{L}_{\text{Prior}}+\lambda_{2}\mathcal{L}_{\text{CLIP}},&\text{for unique images}\\ \lambda_{1}\mathcal{L}_{\text{Prior}}+\lambda_{2}\mathcal{L}_{\text{CLIP}}\\ +\lambda_{3}\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{new}},\mathbf{z}_{\text{ref}}),&\text{for common images}\end{cases}$
where $\mathbf{z}_{\text{adapter, new}}$ and $\mathbf{z}_{\text{adapter, ref}}$ are the adapter-level embeddings for the new and reference subjects, respectively, on the shared common images. We present a breakdown of the high level pipeline loss functions in Appendix I.
@@
The low-level pipeline also follows a similar structure to the high-level one but maps the fMRI data to latent embeddings derived from a custom autoencoder. This autoencoder is trained on downscaled images from the NSD dataset (64x64x3) and reduces them to a latent space of size 1024. The low-level pipeline learns to map to this latent space, and passing these embeddings through a pre-trained decoder produces blurry reconstructions. The loss function for this stage is:
$\mathcal{L}_{\text{new}}=\begin{cases}\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{fMRI, new}},\mathbf{z}_{\text{AE}}),&\text{for unique images}\\ \mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{fMRI, new}},\mathbf{z}_{\text{AE}})\\ +\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{adapt, new}},\mathbf{z}_{\text{adapt, ref}}),&\text{for common images}\end{cases}$
@@
where $\mathbf{z}_{\text{fMRI, new}}$ is the output embedding for the new subject, $\mathbf{z}_{\text{AE}}$ is the AE latent space representation of the image and $\mathbf{z}_{\text{adapter, new}}$ and $\mathbf{z}_{\text{adapter, ref}}$ are the adapter-level embeddings for the new and reference subjects, respectively. We present a breakdown of the low level pipeline loss functions in Appendix J
For the final step, the mapped CLIP embedding is used as a conditioning input to the pre-trained diffusion prior model. The low-level blurry reconstructions serve as an initialization for the diffusion model, and the final, refined reconstructions are then produced. For more architectural details, we refer readers to the MindEye1 (Scotti et al., 2023) and MindEye2 (Scotti et al., 2024) papers.
@@ 3.4 Tests on Complete Data
We first validate our training method using the full dataset. Subject 1 is used as the reference subject, while fine-tuning is performed on Subjects 2, 5, and 7. The results are presented in Table 1. We use the same set of metrics as Scotti et al. (2024) to evaluate our reconstructions. Pixel Correlation, SSIM and 2-way percent correct for the 2nd and 5th layer of AlexNet are considered low level visual metrics. 2-way percent correct for Inception(Szegedy et al., 2016), CLIP, EfficientNet(Tan, 2019) and SwAV(Caron et al., 2020) score are considered high level metrics. When trained to completion, both methods—end-to-end training and adapter alignment—show similar overall performance. However, the MSE loss at the adapter level is significantly lower (by an order of magnitude as shown in Appendix B.2) when using adapter alignment. This allows for more precise post-hoc analysis after extracting embeddings from the common space.
Additionally, we observe that adapter alignment yields better results than regular end-to-end training when the model is trained for fewer epochs, with faster convergence. Both approaches eventually reach similar performance levels, as shown in Figure 3.
Table 1: Quantitative results on training the reconstruction models on all 40 hours of data. Both methods perform similarly across all metrics.
@@
Subj1 Pretraining 0.345 0.346 92.80% 96.88% 94.40% 90.02% 0.692 0.399
Normal FineTune Avg 0.258 0.339 89.64% 95.05% 92.63% 89.64% 0.717 0.427
AAMax FineTune Avg 0.259 0.337 89.20% 95.09% 92.68% 89.49% 0.713 0.423
Subj2 FineTune Normal 0.294 0.345 91.42% 95.92% 93.00% 88.68% 0.721 0.425
Subj2 FineTune AAMax 0.295 0.340 91.52% 96.20% 92.96% 89.87% 0.707 0.409
Subj5 FineTune Normal 0.236 0.335 89.45% 95.32% 93.98% 91.63% 0.701 0.419
Subj5 FineTune AAMax 0.240 0.332 88.60% 95.16% 94.04% 90.71% 0.698 0.413
Subj7 FineTune Normal 0.244 0.339 88.06% 93.91% 90.92% 88.61% 0.731 0.438
Subj7 FineTune AAMax 0.242 0.340 87.48% 93.91% 91.05% 87.89% 0.734 0.448
Figure 3: Comparing finetuning reconstruction performance over time with all the data. From left to right: (a) High level metrics over time for Subject 1. Both methods perform similarly but the normal method takes about 20 epochs to match AAmax’s numbers (b) Example reconstruction over time. The top row is the normal pipeline and the bottom row is the Adapter alignment pipeline.
@@ 3.5 Tests on Limited Data
A key objective of multi-subject pipelines is to achieve satisfactory models of new subjects with limited data. Acquiring fMRI data is costly, and reducing the data requirements for modeling a subject’s brain makes fMRI reconstruction more accessible. MindEye2 demonstrated that fine-tuning a new subject with as little as 1 hour of data is possible, though the resulting reconstructions leave room for improvement.
In our experiments, we evaluate the fine-tuning of a new subject in three scenarios: 1 hour, 2 hours, and 4 hours of training data. The NSD dataset provides 40 hours of data per subject, all of which is used for the full data tests in Section 3.4. For this experiment, we designate Subject 5 as the reference subject and fine-tune Subject 1 using the shared images. We compare the performance of normal end-to-end training and Adapter Alignment (AA), with results presented in Table 11.
Adapter Alignment consistently outperforms the end-to-end approach, showing significant improvements even after just one epoch of training. The performance gap widens as the amount of fine-tuning data is reduced. Due to the lightweight nature of Adapter Alignment, which involves training on just 250 images using a single layer, the time required to align the adapter is minimal compared to full end-to-end training. We present the training times for both methods in Appendix Section G. In fact, even after 160 epochs of normal training, the model’s performance fails to match the Epoch 1 results of Adapter Alignment. Figure 4 illustrates the reconstruction performance across epochs for both methods.
Figure 4: Comparing finetuning reconstruction performance over time with limited data. From left to right: (a) Finetuning Subject 1 with 250 images or 1 hour of data (b) Finetuning Subject 1 with 500 images or 2 hours of data (c) Finetuning Subject 1 with 1000 images or 3 hours of data. We observe that adapter alignment (AAMax) not only outperforms normal end-to-end training, but the performance at Epoch 1 is better than the normal method’s performance at Epoch 160.
Table 2: Quantitative results on fine tuning a new subject with limited data. FT=Fine tuning. The first section shows the result of fine tuning with 250 images after 1 epoch and after 160 epochs using normal end-to-end training vs AAMax. The second section compares performance on 500 images and the third section compares performance on 1000 images. In all cases AAMax, significantly outperforms normal training. The greyed out low level metrics are not optimized for in these experiments.
@@
Subj1 FT NS 250 Normal E1 0.031 0.234 58.31% 62.56% 58.51% 58.71% 0.947 0.611
Subj1 FT NS 250 AAMAX E1 0.095 0.254 77.20% 85.98% 78.80% 79.39% 0.833 0.503
Subj1 FT NS 250 Normal E160 0.094 0.239 71.54% 80.10% 73.11% 74.02% 0.875 0.537
Subj1 FT NS 250 AAMAX E160 0.105 0.259 76.77% 86.20% 80.08% 80.14% 0.828 0.496
Subj1 FT NS 500 Normal E1 0.049 0.242 62.41% 68.06% 62.64% 64.80% 0.925 0.583
Subj1 FT NS 500 AAMAX E1 0.112 0.275 79.46% 88.29% 83.28% 82.61% 0.803 0.485
Subj1 FT NS 500 Normal E160 0.110 0.270 75.32% 84.62% 76.66% 78.55% 0.845 0.516
Subj1 FT NS 500 AAMAX E160 0.114 0.285 80.00% 88.50% 83.19% 83.05% 0.799 0.480
Subj1 FT NS 1000 Normal E1 0.079 0.241 65.84% 73.04% 64.19% 68.10% 0.914 0.574
Subj1 FT NS 1000 AAMAX E1 0.129 0.265 80.83% 90.93% 86.66% 86.17% 0.767 0.460
Subj1 FT NS 1000 Normal E160 0.124 0.253 77.86% 89.18% 83.10% 83.26% 0.801 0.480
Subj1 FT NS 1000 AAMAX E160 0.127 0.273 81.73% 91.29% 86.23% 86.26% 0.763 0.453
4 Best Candidate Image Selection
Here, we present a selection algorithm that significantly reduces the amount of data required to train a model for a new subject. Later, we carry out an empirical analysis of the method.
4.1 Greedy Algorithm
In the context of fMRI-based visual reconstruction, we introduce a greedy algorithm that selects the most representative subset of images from the common space. We begin with an embedding of the common set of images into a representation space and choose the first $d$ dimensions corresponding to the principal singular vectors. For each dimension $j$, we partition the range into $B_{j}$ bins that is determined by the ratio of singular value $\lambda_{j}$ to the largest singular value $\lambda_{1}$:
Each image (and its fMRI) maps to exactly one bin in each dimension. Our goal is to find the smallest subset of images that covers all bins in each dimension. This subset is critical in selecting the most representative images for a new subject, ensuring that the selected images capture the key activity patterns that generalize across subjects. The problem as stated is NP-hard. However, it is submodular and a greedy heuristic such as given below achieves an $(1-1/e)$ approximation ratio.
After acquiring the selected images, we train a ridge regression layer on this subset, which then serves as an adapter for the remainder of the pipeline. With only a minimal number of additional end-to-end training epochs, our approach achieves comparable performance while relying on a substantially smaller training set. Section F in the Appendix provides additional details about the algorithm, while Section E contains the proof of the submodularity property.
4.2 Results
We achieve better results on high-level metrics compared to the baseline while reducing the required data by 50%. Figure 5 (a) illustrates how Inception and CLIP distances fluctuate as a function of the amount of training data for subject 2. Figure 5 (b) presents the model’s performance when 150 images are selected for the training set, while varying the number of principal eigenvectors onto which the embeddings are projected. Differences in the performance for different numbers of dimensions are likely due to the nature of the binning process, as the number of bins created along the eigenvectors provides better coverage. Further analysis of this behavior presents an interesting direction for future work, especially considering that, as Appendix H demonstrates, the optimal number of eigenvectors varies across subjects. Our method highlights the critical role of strategic data selection in achieving superior performance. For additional details on the experimental performance, please refer to section H in the Appendix.
Figure 5: Comparison of the performance of the image selection strategy. Both figures display the Inception and CLIP distance metrics for subject 2. (a) Performance with varying training set sizes, where the red and black lines (labeled as Inception RF and CLIP RF) at the 250-point mark indicate the model’s performance without a data selection strategy. (b) Performance with varying numbers of principal eigenvectors.
Finally, the most significant performance gains emerge when image selection is combined with adaptive alignment. As shown in Table 3, this integrated approach outperforms conventional finetuning.
Table 3: Performance metrics when fine-tuning for Subject 2. Regular FT stands for regular fine-tuning (row 1). AA stands for adapter alignment (row 2). AA + IS stands for adapter alignment with image selection (row 3).

### v2

@@ 4.3 Reconstruction Pipeline
In general, fMRI-to-image reconstruction pipelines follow a common structure: brain signals are mapped into an intermediate embedding space (semantic or virtual), which is then decoded into an image using a pre-trained generative model such as a diffusion decoder. We discuss multiple previous approaches and their pipelines in Appendix 2. Since Adapter Alignment (AA) is agnostic to architectural choices, we adopt the pipeline used in MindEye1 for our main experiments and refer readers to MindEye1 (Scotti et al., 2023) and MindEye2 (Scotti et al., 2024) for architectural details. We also validate our method on MindEye2’s architecture (Section 5.4 and Appendix A) and across varying component choices and shared space dimensionalities (Appendix H), showing consistent improvements and compatibility.
@@
In summary, flattened fMRI voxels from the “nsdgeneral” region are passed through a subject-specific adapter (a single linear layer with GELU non-linearity) into a shared representation space. This shared representation is then processed by a residual MLP. The resulting output is mapped into the CLIP (Radford et al., 2021) space via a pre-trained diffusion prior (Ramesh et al., 2022). This serves as a conditioning signal for a diffusion model to produce the final reconstruction.
We detail the training losses used in this pipeline for both the high-level (semantic) and low-level (visual) branches below, with full equations provided in Appendices M and N. The overall loss function for the high-level pipeline is:
@@
$\mathcal{L}_{\text{new}}=\begin{cases}\lambda_{1}\mathcal{L}_{\text{Prior}}+\lambda_{2}\mathcal{L}_{\text{CLIP}},&\text{for unique images}\\ \lambda_{1}\mathcal{L}_{\text{Prior}}+\lambda_{2}\mathcal{L}_{\text{CLIP}}+\lambda_{3}\mathcal{L}_{\text{MSE}}(\mathbf{z}_{\text{new}},\mathbf{z}_{\text{ref}}),&\text{for common images}\end{cases}$
@@
The low-level pipeline also follows a similar structure to the high-level pipeline but maps the fMRI data to latent embeddings derived from a custom autoencoder. This autoencoder is trained on downscaled images from the NSD dataset (64x64x3) and reduces them to a latent space of size 1024. The low-level pipeline learns to map to this latent space; passing these embeddings through a pre-trained decoder produces blurry reconstructions. The loss function for this stage is:
[…]
@@
where $\mathbf{z}_{\text{fMRI, new}}$ is the output embedding for the new subject, $\mathbf{z}_{\text{AE}}$ is the AE latent space representation of the image and $\mathbf{z}_{\text{adapter, new}}$ and $\mathbf{z}_{\text{adapter, ref}}$ are the adapter-level embeddings for the new and reference subjects, respectively.
After mapping brain signals to the CLIP space, the mappings are fed as conditioning to a frozen Diffusion Model (Xu et al., 2023) to generate final reconstructions.
4.4 Greedy Algorithm for Best Candidate Image Selection
In the context of fMRI-based visual reconstruction, we introduce a greedy algorithm that selects the most representative subset of images from the common space. We first decompose (SVD) the adapter’s weight matrix of the reference subject. Then we project the embeddings of the shared set of 1000 images onto the $d$ principal dimensions corresponding to the principal singular vectors. For each dimension $j$, we partition the range into $B_{j}$ bins that is determined by the ratio of singular value $\lambda_{j}$ to the largest singular value $\lambda_{1}$:
Each image (and its fMRI) maps to exactly 1 bin in each dimension and our goal is to find the smallest subset of images that covers all bins in each dimension. This subset is critical in selecting the most representative images for a new subject, ensuring that the selected images capture the key activity patterns that generalize across subjects. The problem as stated is NP-hard. However, it is submodular and a greedy heuristic such as given below achieves an (11/e)(1-1/e) approximation ratio.
After applying the image selection algorithm, the chosen subset simply replaces random sampling for fine-tuning. The training procedure itself is unchanged: the selected images can be used with either baseline fine-tuning or AAMax. In this way, the algorithm serves purely as a strategy to identify the most representative subset of images, reducing the data requirement without altering the reconstruction pipeline. Formal proofs for the algorithm are presented in Appendix K.
5 Results
We now evaluate the performance of our proposed framework across a range of experimental conditions. We begin with complete data experiments, where AAMax is compared against baseline training. We then turn to limited data settings, where we report results from both AAMax and AAMax combined with image selection in a unified comparison. We also analyze how AAMax scales with increasing amounts of training data, highlighting that its advantage diminishes as larger datasets allow even unaligned models to generalize. Finally, we present experiments showing that AAMax generalizes across architectures using MindEye1 and MindEye2 (Scotti et al., 2024) architectures.
@@ 5.1 Tests on Complete Data
We first validate our training method using the full dataset. Subject 1 is used as the reference subject, while fine-tuning is performed on Subjects 2, 5, and 7. The results averaged over all 3 subjects are presented in Table 1. AA is agnostic to the choice of reference subject but we found that having a good reference subject improves fine-tuning performance (Appendix C). We use the same set of metrics as Scotti et al. (2024) to evaluate our reconstructions. Pixel Correlation, SSIM and 2-way percent correct for the 2nd and 5th layer of AlexNet are considered low level metrics. 2-way percent correct for Inception(Szegedy et al., 2016), CLIP, EfficientNet(Tan, 2019) and SwAV(Caron et al., 2020) score are considered high level metrics. When trained to completion, both baseline fine-tuning and AAMax show similar overall performance. However, the MSE loss at the adapter level is significantly lower (by an order of magnitude as shown in Appendix I.2) when using adapter alignment. This allows for more precise post-hoc analysis after extracting embeddings from the common space.
Table 1: Quantitative results on training the reconstruction models on all 40 hours of data. The fine-tuning results are averaged over subjects 2,5 and 7. AAMax and baseline fine-tuning perform similarly across all metrics.
@@
Subject 1 Pre-train 0.345 0.346 92.80% 96.88% 94.40% 90.02% 0.692 0.399
Baseline Fine-Tune AVG 0.258 0.339 89.64% 95.05% 92.63% 89.64% 0.717 0.427
AAMax Fine-Tune AVG 0.259 0.337 89.20% 95.09% 92.68% 89.49% 0.713 0.423
@@ 5.2 Tests on Limited Data
A key objective of multi-subject pipelines is to achieve accurate models of new subjects with limited data. Because acquiring fMRI data is costly, reducing the data requirements makes reconstruction pipelines more practical and broadly applicable.
In our experiments, we designate Subject 1 as the reference and fine-tune Subjects 2, 5, and 7 using the shared images. For consistency, we use the same reference subject checkpoint across all experiments and report averaged results over three runs with fixed random splits. Table 2 presents the unified results for fine-tuning with 1 hour of data (250 images), comparing baseline training and AAMax with both random sampling and image selection (IS). Across all subjects, AAMax significantly outperforms the baseline, while IS consistently improves performance regardless of the fine-tuning strategy. As expected, the strongest results are obtained when combining AAMax with IS, showing that the two approaches are complementary. For IS, we select images based on the top 20 eigenvector dimensions, with ablations on this choice provided in Appendix D. We run experiments with scaling the amount of training data in the following section and additional ablations with 2 and 4 hours of data are included in Appendix F, where the same trend holds. We also run experiments on convergence in Appendix F and find that AA followed by just 1 epoch of End-to-End fine-tuning is good enough to outperform the baseline.
We further investigate the efficiency of our approach by testing how far data requirements can be reduced while maintaining reconstruction performance. In this setup, Subjects 2 and 5 were fine-tuned from Subject 1 using 100, 150, and 200 training images. At 100 images, both AAMax and AAMax+IS already match the baseline trained with 1 hour of data, and at 150 images they clearly surpass it. As shown in Figure 3, AAMax+IS consistently achieves the strongest results, though even AAMax alone outperforms the baseline in these low-data regimes. These findings highlight the robustness of adapter alignment and demonstrate that effective reconstructions can be achieved with as little as 40% of the data previously considered necessary.
Table 2: Limited data fine-tuning results (1 hour, 250 images) using Subject 1 as the reference and Subjects 2, 5, and 7 as fine-tuning subjects. Each subject is evaluated under four settings: Baseline with random sampling, AAMax with random sampling, Baseline with image selection (IS), and AAMax with IS. Results are averaged over three runs with fixed random splits. AAMax consistently outperforms the baseline, IS provides additional gains in both training regimes, and the best performance is achieved by combining AAMax with IS. Pixel-level metrics are shown in gray since they are not optimized in these runs.
@@
Subject 2 FT 1HR Baseline 0.078 0.264 75.33% 84.18% 74.15% 73.66% 0.862 0.524
Subject 2 FT Selected 1HR Baseline 0.097 0.288 78.78% 85.78% 77.13% 76.17% 0.842 0.513
Subject 2 FT 1HR AAMax 0.112 0.259 78.75% 88.04% 78.99% 78.28% 0.824 0.513
Subject 2 FT Selected 1HR AAMax 0.102 0.267 81.41% 89.55% 80.43% 80.33% 0.823 0.496
Subject 5 FT 1HR Baseline 0.082 0.267 75.05% 83.28% 76.04% 76.40% 0.855 0.526
Subject 5 FT Selected 1HR Baseline 0.098 0.286 77.07% 86.31% 78.09% 79.55% 0.827 0.499
Subject 5 FT 1HR AAMax 0.105 0.278 79.90% 88.68% 80.63% 82.20% 0.808 0.493
Subject 5 FT Selected 1HR AAMax 0.117 0.285 81.05% 89.88% 81.92% 82.47% 0.799 0.480
Subject 7 FT 1HR Baseline 0.079 0.252 72.98% 79.98% 71.57% 72.07% 0.871 0.539
Subject 7 FT Selected 1HR Baseline 0.099 0.286 78.41% 84.77% 74.91% 75.79% 0.844 0.515
Subject 7 FT 1HR AAMax 0.104 0.285 80.04% 87.47% 79.17% 78.93% 0.826 0.509
Subject 7 FT Selected 1HR AAMax 0.123 0.298 81.24% 88.34% 79.94% 79.81% 0.820 0.495
Figure 3: Pushing the limits of data efficiency with AAMax+Image Selection. Left: Subject 2 fine-tuned from Subject 1. Right: Subject 5 fine-tuned from Subject 1. In both cases, AAMax and AAMax+Image Selection outperform the baseline trained with 250 images, even when using only 150 images (a 40% reduction) and match performance with 100 images (a 60% reduction).
5.3 Scaling with Training Data
We further examine how the benefits of Adapter Alignment scale with training data size. Using Subject 2 as the fine-tuning subject, we progressively increase the training set from 250 to 500, 1000, and 4000 images, before extending to the full dataset. Up to 1000 images, only common images are used for alignment and fine-tuning; for 4000 images, alignment is still performed on the 1000 shared images, while the additional 3000 images are used only for fine-tuning. As shown in Figure 4, AAMax achieves substantially higher performance than baseline fine-tuning in low-data regimes, with the gap narrowing as training data increases. At full data, both approaches converge, consistent with prior multi-subject pipelines such as MindEye2, where large amounts of subject-specific data allow the model to generalize even without explicit alignment.
Figure 4: Performance of AAMax compared to baseline fine-tuning as a function of training data size. Left: AlexNet-5 2-way percent correct. Middle: Inception 2-way percent correct. Right: CLIP 2-way percent correct. Across all three metrics, AAMax achieves higher accuracy in limited-data regimes, with the advantage diminishing as more data becomes available.
5.4 Generalizing To Different Architectures
Adapter Alignment is inherently architecture-agnostic: any reconstruction model can be extended with subject-specific adapters, enabling alignment to a common space. For efficiency and comparability with prior work, all of our main experiments are conducted using a slightly modified MindEye1 pipeline. To demonstrate the architecture-agnostic nature of Adapter Alignment, we additionally evaluate a limited-data setting using 1 hour of training data to fine-tune Subject 2. We compare baseline fine-tuning and AAMax under both the MindEye1 and MindEye2 architectures, and observe consistent improvements with AAMax across both models. Additional experiments using MindEye2 are presented in Appendix A.
Table 3: Evaluating Adapter Alignment across architectures. Subject 1 is trained as the reference, and Subject 2 is fine-tuned with 1 hour of training data. Results are shown for MindEye1 and MindEye2 under the baseline and AAMax fine-tuning. These results demonstrate that AA is architecture-agnostic. Moreover, AA scales with the underlying architecture—stronger models such as MindEye2 achieve higher absolute performance, with their gains further enhanced by $A$.